Please show formulas and work. No excel formulas please.
Problem 1 You take out an amortized loan for $10,000. The loan is to be paid in equal installments at the end of each of the next 5 years. The interest rate is 8%. Construct an amortization schedule. Problem 2: A. Calculate the PV of $100 due in 5 years compounded daily at 12%. B. Calculate the FV of $1000 due in 3 years at 6% compounded quarterly. C. Calculate the FVA of $300 due at the end of each of the next 5 years at 496. D. Calculate the PVA of $300 due at the end of each of the next 5 years at 4%. Problem 3: Compute the EAR of 10% compounded daily. Previous Next search 0
Answer #1
1) Amortization Schedule for 5 Years:
Year Beginning Loan balance Interest Expense Annual Payment Reduction in Loan balance Ending Loan Balance
a b=a*8% c d=c-b e=a-d
1 $ 10,000.00 $     800.00 $   2,504.56 $ 1,704.56 $ 8,295.44
2 $   8,295.44 $     663.63 $   2,504.56 $ 1,840.93 $ 6,454.51
3 $   6,454.51 $     516.36 $   2,504.56 $ 1,988.20 $ 4,466.30
4 $   4,466.30 $     357.30 $   2,504.56 $ 2,147.26 $ 2,319.04
5 $   2,319.04 $     185.52 $   2,504.56 $ 2,319.04 $         0.00
Total $ 2,522.82 $ 12,522.82
Working:
# 1 Present value of annuity of 1 = (1-(1+i)^-n)/i where,
= (1-(1+0.08)^-5)/0.08 i 8%
= 3.99271004 n 5
# 2 Annual Payment = Loan Amount/Present value of annuity of 1
= $       10,000 / 3.99271
= $   2,504.56